2xy^2dx=e^x(dy-ydx) solve the above diffrantial eqation 

Grade:Select Grade

1 Answers

jagdish singh singh
8 years ago
\hspace{-0.6 cm }$Given $\bf{2xy^2dx=e^x(dy-ydx)\Rightarrow 2xdx = \frac{e^x(dy-ydx)}{y^2}}$\\\\\\So we get $\bf{-2xdx = \frac{ye^xdx-e^xdy}{y^2} =d\left( \frac{e^x}{y}\right)\Rightarrow d\left( \frac{e^x}{y}\right)=-2xdx}$\\\\\\ So we get $\bf{\frac{e^x}{y}=-x^2+\mathcal{C} \Rightarrow y=\frac{e^x}{-x^2+\mathcal{C}}.}$